# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

### v1

@@ 3.3 The optimal perturbation in the pilot stage
Solving (23) gives $\gamma=-1/10$ and $3V_{K}\sigma^{2}(\theta_{0})=4H_{K}^{2}$. From Theorem 2, we can calculate $V_{K}\neq 0$ easily. Usually $\sigma^{2}(\theta_{0})>0$, so $H_{K}\neq 0$, which implies $D\neq 0$. This gives the rationality of $\alpha^{(5)}(\theta_{0})\neq 0$ in Assumption 3.
On the other hand, if $c_{1},...,c_{K}$ are randomly generated from $\mathcal{P}_{0}$, the optimization problem comes from the combination of (19) and (3.2.1):
@@
Solving (24) also yields $\gamma=-1/10$. Thus, whether $c_{1},...,c_{K}$ are fixed or randomly generated, the optimal selection of $h_{k}$’s in the pilot stage is of order $O\left(n_{b}^{\gamma}\right)$ with $\gamma=-1/10$. It is noteworthy that the optimal perturbation in the pilot stage differs from the optimal one for the CFD problem shown in (8).
A more exact determination of $\{c_{1},...,c_{K}\}$ or $\mathcal{P}_{0}$ requires knowledge of $D$, which is more challenging than the estimation of $\alpha^{\prime}(\theta_{0})$. In this paper, we treat $\{c_{1},...,c_{K}\}$ and $\mathcal{P}_{0}$ as the hyperparameters whose selection is beyond the scope of our discussion.
4 Sample-recycling for Finite Difference Output
In Section 3, we have provided the estimators $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ and shown that both of them are consistent. In this section, we use $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ to set the estimated optimal perturbation by (8) and construct the estimators for $\alpha^{\prime}(\theta_{0})$.
The straightforward method is neglecting the $Kn_{b}$ pilot samples, i.e., setting
@@ 4 Sample-recycling for Finite Difference Output
where $n_{2}=n-Kn_{b}$. Then, $n_{2}$ sample pairs are used to construct $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ by (3) as the estimator of $\alpha^{\prime}(\theta_{0})$. From (9), given $\widehat{h}_{n_{2}}$, the expectation and variance of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ are respectively
@@
According to $\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\mathbb{E}\left[\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]$ and $\mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\mathrm{Var}\left[\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]+\mathbb{E}\left[\mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]$, we have [^5] To be rigorous, (27) requires $o_{p}(\cdot)$ and $o_{p}^{2}(\cdot)$ in (26) to be uniformly integrable. Throughout the paper, we assume this condition implicitly when necessary to prevent excessive technical complexity. The rationality of the leading terms in (27) is referred to in Appendix B.1.
@@
Li and Lam (2020) presents a similar result to (27). However, their approach to constructing $\widehat{h}_{n_{2}}$ differs from ours. For more details, interested readers may refer to Li and Lam (2020).
From (27), the MSE of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ is $O\left(n_{2}^{-2/3}\right)$. When $n_{b}=o(n)$, the leading term of the MSE of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ matches that of $\widehat{\Delta}_{n,h^{*}}$, where $h^{*}$ is the optimal perturbation as defined in (8). However, this result may not hold in some practical scenarios where the sample size is not large enough.
When $n_{b}\neq o(n)$ and $Kn_{b}=rn$, where $r$ is the proportion of the pilot samples and $10<r<$ 1, the MSE of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ is
$3\left(\dfrac{B^{2}\bm{\beta}_{v}^{2}}{16(1-r)^{2}n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).$
Evidently, although this MSE is also of order $n^{2/3}$, it is larger than the optimal one in (9).
Therefore, the estimator $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ lacks efficiency because of its construction using only $n_{2}(<n)$ sample pairs. An intuitive way of improving the efficiency is to recycling the pilot samples in constructing the estimator. In the next section, we combine a pilot samples and the $n_{2}$ sample pairs to establish an efficient estimator.
4.1 Constructing the DSR-CFD estimator
In the following, we construct the double sample-recycling CFD (DSR-CFD) estimator through the remain $n_{2}$ sample pairs as well as the pilot sample pairs. We have to consider two questions: (1) how to generate the $n_{2}$ sample pairs; and (2) how to recycle the pilot sample pairs.
To answer the first question, note that we hope to utilize the whole $n$ sample pairs, including the pilot sample pairs and the $n_{2}$ sample pairs. Therefore, the desired CFD estimator should be constructed based on the perturbation $\widehat{h}_{n}$, where
@@ 4.1 Constructing the DSR-CFD estimator
Note that $\widehat{h}_{n}$ is derived from the entire sample-pair set of size $n$, whereas $\widehat{h}_{n_{2}}$, as defined in (25), is based on the subset of size $n_{2}$. Next, $n_{2}$ sample pairs are generated at $\widehat{h}_{n}$, and we can calculate $\left\{\Delta_{i}\left(\widehat{h}_{n}\right),i=1,...,n_{2}\right\}$ based on (2). When no confusion arises, the subscript $i$ of $\Delta_{i}\left(\widehat{h}_{n}\right)$ is omitted to denote a generic $\Delta\left(\widehat{h}_{n}\right)$.
@@
This is a weighted average of an estimator at $\widehat{h}_{n}$ constructed through the $n_{2}$ samples, and the one by the transformation from the pilot samples. It is reasonable to expect that under certain mild conditions, such as ensuring that the covariance of two terms in (30) converges faster than the variance of either term, the bias, variance and MSE of (30) all achieve their optimal values as in (9).
However, expectations and variances of $\Delta(h_{k})$ and $\Delta\left(\widehat{h}_{n}\right)$ in (29) are all unknown. Therefore, we have to substitute them in (30) by their estimators. In light of results in Section 3, we choose estimators as shown in Table 2.
@@
In Table 2, the estimators are all constructed through solutions to regression problems (14) and (16). Hence, they only depend on the pilot samples, and are independent with the remaining $n_{2}$ samples conditional on $\widehat{h}_{n}$, facilitating the theoretical analysis of the subsequently proposed estimator.
@@
$\Delta_{j}^{t}(h_{k})=\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}.$ (31)
Then, the DSR-CFD estimator, i.e., the estimated version of (30), is established as follows.
$\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+\sum_{k=1}^{K}\sum_{j=1}^{n_{b}}\Delta_{j}^{t}(h_{k})\right).$ (32)
Next, we analyze the theoretical properties of $\widehat{\Delta}_{n,\widehat{h}_{n}}$. The results are summarized in the following theorem and proposition. To simplify the proof, we consider the non-random hyperparameters $c_{1},...,c_{K}$ such that when $n_{b}$ is given, $h_{1},...,h_{K}$ are also given. The proof is similar but more tedious when considering $\mathcal{P}_{0}$ as the hyperparameter, i.e., $c_{1},...,c_{K}$ are random.
@@
$\mathbb{E}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\alpha^{\prime}(\theta_{0})+\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}n^{-1/3}+o\left(n^{-1/3}\right),$ (33)
$\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}\left[\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}-K\right]n^{-2/3}+o\left(n^{-2/3}\right),$ (34)
$\mathrm{MSE}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $3\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3}+\left\{\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}\left(\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}+{\rm{sign}}(B)\sigma(\theta_{0})\right)\right.$
$+\left.\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}\left[\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}-K\right]\right\}n^{-2/3}+o\left(n^{-2/3}\right),$ (35)
where ${\bm{c}}=[|c_{1}|,...,|c_{K}|]^{\top}$, ${\bm{c}^{4}}=[c_{1}^{4},...,c_{K}^{4}]^{\top}$, ${\bm{P}}=\bm{I}-\bm{X}_{e}(\bm{X}_{e}^{\top}\bm{X}_{e})^{-1}\bm{X}_{e}^{\top}$, $\bm{X}_{e}$ is defined as in (15), ${\rm{Diag}}(\bm{c^{-1}})={\rm{Diag}}\left(\dfrac{1}{|c_{1}|},...,\dfrac{1}{|c_{K}|}\right)$, ${\rm{sign}}(B)$ is the sign of $B$, and $||{\bm{v}}||_{2}=\left(\sum_{k=1}^{K}v_{k}^{2}\right)^{1/2}$ for any ${\bm{v}}\in\mathbb{R}^{K}$.
The proof of Theorem 4 is provided in Appendix B.2. Comparing (33), (34) and (35) with (9), the changes of bias, variance, and MSE in the DSR-CFD method compared to the optimal CFD method are evident, which depend on ${\bm{c^{\top}Pc^{4}}}$ and $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}$. In Appendix B.3, we indicate that these changes are typically manageable.
Even more attractive is the fact that if $B$ and $D{\bm{c^{\top}Pc^{4}}}$ are of opposite signs, the squared bias of $\widehat{\Delta}_{n,\widehat{h}_{n}}$ will be smaller than that of the optimal CFD estimator (see (9)). In addition, when $c_{1},...,c_{K}$ are chosen such that $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$, the variance will be reduced. Proposition 5, a corollary of Theorem 4, further elaborates on these aspects.
@@
      If $-1-{\rm{sign}}(B)\leq\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}D{\bm{c^{\top}Pc^{4}}}\leq 1-{\rm{sign}}(B)$, then
@@
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|\leq\left(1+\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}\left|D{\bm{c^{\top}Pc^{4}}}\right|\right)\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right).$
@@
These results in Proposition 5 can be easily derived by the same technique as in Theorem 4 and the proof is provided in Appendix B.4. Proposition 5 emphasizes that the DSR-CFD estimator possesses a reduced variance and in some cases, a reduced bias, compared to the optimal CFD estimator, and it is applicable even if $r=1$. Specifically, Proposition 5 guides the following conclusions:
@@
    The first result indicates that the DSR-CFD method regains the CFD method when $n_{b}=o(n)$ (refer to (9) and (36)). This aligns with intuition since we allocate a majority of the samples to gradient estimation rather than constant estimation.
@@
    The second result in Proposition 5 suggests that if $n_{b}=O(n)$, then the absolute value of the bias of the DSR-CFD estimator converges at the same rate as that of the optimal CFD estimator, or it may even be reduced compared to the latter. In particular, 2.i in Proposition 5 demonstrates that the former is smaller than the latter. On the other hand, in Appendix B.3, we show that ${\bm{c^{\top}Pc^{4}}}$ is typically controllable. Therefore, 2.ii in Proposition 5 illustrates that even if the absolute value of the bias of the DSR-CFD estimator is not reduced, its increment is moderate and will not blow up.
@@
    The third result in Proposition 5 also demonstrates that if $n_{b}=O(n)$ and $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$, the variance of the DSR-CFD estimator is reduced compared to that of the optimal CFD estimator. It is worth mentioning that the condition $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ is a highly sufficient condition of $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$, and the latter leads to a reduced variance (see Theorem 4).
@@
    An appealing observation from Proposition 5 is the robustness of the DSR-CFD method w.r.t. $r$, in the sense that our proposed DSR-CFD estimator exhibits favorable theoretical performance for both $r=0$ and $r>0$. In particular, when $r=1$, i.e., indicating that the entire simulation budget is allocated to the pilot phase, the DSR-CFD estimator remains efficient. This theoretical insight serves as a guide, suggesting that when faced with limited simulation resources, allocating the entire budget to generate pilot samples, and using the double sample-recycling technique can lead to an efficient finite difference estimator.
Remark 2.
The efficiency of the DSR-CFD estimator is attributed to the expectation and variance of (29) being identical to those of the optimal CFD estimator. However, by substituting $\mathbb{E}[\Delta(h_{k})]$, $\mathbb{E}\left[\Delta\left(\widehat{h}_{n}\right)\right]$, $\mathrm{Var}[\Delta(h_{k})]$ and $\mathrm{Var}\left[\Delta\left(\widehat{h}_{n}\right)\right]$ using their regression estimators, respectively, so we introduce the intrinsic correlation and this may reduce the variance as well as the bias. In fact, the reason for variance reduction and possible bias reduction lies in the substitution of $\mathbb{E}[\Delta(h_{k})],k=1,...,K$. If we impractically assume that $\mathbb{E}\Delta(h_{k}),k=1,...,K$ are all known precisely and denote $\Delta_{j}^{t}(h_{k})$ by
$\Delta_{j}^{t}(h_{k})=\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e},$
then the new DSR-CFD estimator, denoted by $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$, fully recovers the result of the optimal CFD estimator. In Appendix B.5, we provide these properties. After substituting $\mathbb{E}[\Delta(h_{k})],k=1,...,K$ with their regression estimators $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e},k=1,...,K$, two scenarios may occur:

### v4

@@ 3.3 The Optimal Perturbation for Pilot Samples
Solving (23) gives $\gamma=-1/10$. On the other hand, if $c_{1},...,c_{K}$ are generated from $\mathcal{P}_{0}$ randomly, the optimization problem comes from the combination of (19) and (3.2.1):
@@
Solving (24) also gives $\gamma=-1/10$. That is, whether $c_{1},...,c_{K}$ are fixed or not, the optimal selection of $h_{k}$’s are of order $n_{b}^{\gamma}$ with $\gamma=-1/10$.
A more exact setting about $\{c_{1},...,c_{K}\}$ or $\mathcal{P}_{0}$ necessitates knowledge of $D$, which is more challenging than the estimation of $\alpha^{\prime}(\theta_{0})$. In this paper, we treat $\{c_{1},...,c_{K}\}$ and $\mathcal{P}_{0}$ as the hyperparameters whose selection is beyond the scope of our discussion.
4 The Correlation-induced FD Estimator
In this section, we use estimators $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ to estimate the optimal perturbation by (8) and construct the correlation-induced estimator for $\alpha^{\prime}(\theta_{0})$.
The straightforward method for estimating the optimal perturbation is neglecting the $Kn_{b}$ pilot sample pairs, i.e., setting
@@ 4 The Correlation-induced FD Estimator
where $n_{2}=n-Kn_{b}$. Then, $n_{2}$ sample pairs are used to construct $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ by (3) as the estimator of $\alpha^{\prime}(\theta_{0})$. It is noteworthy that this estimator is similar to the one constructed in Li and Lam (2020), with the primary difference being the estimations of $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$. From (9), given $\widehat{h}_{n_{2}}$, the expectation and variance of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ are respectively
@@
According to $\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\mathbb{E}\left[\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]$ and $\mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\mathrm{Var}\left[\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]+\mathbb{E}\left[\mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]$, we have
@@
Because $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ is not the one we study in this paper, the discussion and proof of (27) are provided in Appendix B.1.
It is worth noting that $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ only uses $n_{2}$ $(<n)$ sample pairs. Although the MSE shown in (27) achieves the theoretical optimality as shown in (9) when $n_{b}=o(n)$, neglecting the pilot samples results in reduced efficiency when the sample size is small.
4.1 Constructing the Estimator by Inducing Correlated Samples
To avoid neglecting pilot samples in the construction of CFD estimators, we attempt to combine the pilot samples with the additional $n_{2}$ sample pairs. In fact, as will be indicated in Section 4.2, theoretical analyses suggest that $n_{2}=0$ is still feasible. Next, we must address two questions: (1) how to generate the $n_{2}$ sample pairs; and (2) how to reuse the pilot samples.
To answer the first question, because we hope to utilize the whole $n$ sample pairs, the desired CFD estimator should be constructed based on the perturbation $\widehat{h}_{n}$, where
@@ 4.1 Constructing the Estimator by Inducing Correlated Samples
Note that $\widehat{h}_{n}$ is derived from the entire sample-pair set of size $n$, whereas $\widehat{h}_{n_{2}}$, as defined in (25), is based on the subset of size $n_{2}$. Next, the additional $n_{2}$ sample pairs are generated at $\widehat{h}_{n}$, and we can calculate $\left\{\Delta_{i}\left(\widehat{h}_{n}\right),i=1,...,n_{2}\right\}$ based on (2). When no confusion arises, the subscript $i$ of $\Delta_{i}\left(\widehat{h}_{n}\right)$ is omitted to denote a generic $\Delta\left(\widehat{h}_{n}\right)$.
@@
This is a weighted average of an estimator at $\widehat{h}_{n}$ constructed through the $n_{2}$ sample pairs and the one by the transformation from the pilot samples. It is reasonable to expect that under mild conditions, such as ensuring that the covariance of two terms in (30) converges faster than the variance of either term, the bias, variance and MSE of (30) all achieve their optimal values as in (9).
However, expectations and variances of $\Delta(h_{k})$ and $\Delta\left(\widehat{h}_{n}\right)$ in (29) are both unknown. Therefore, we have to substitute them in (30) by their estimators. In light of results in Section 3, we choose estimators as shown in Table 2.
@@
Specifically, the estimators are all constructed through solutions to regression problems (14) and (16).
@@
$\Delta_{j}^{cor}(h_{k})=\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}.$ (31)
Compared with (29), the transformation induces correlation to $\Delta_{j}^{cor}(h_{k})$ for $j=1,...,n_{b}$, and so we use the superscript “$cor$” to highlight the correlation. Then, the correlation-induced CFD (Cor-CFD) estimator, i.e., the estimated version of (30), is established as follows.
$\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+\sum_{k=1}^{K}\sum_{j=1}^{n_{b}}\Delta_{j}^{cor}(h_{k})\right).$ (32)
4.2 Asymptotic Analyses of the Cor-CFD Estimator
The results for the Cor-CFD estimator $\widehat{\Delta}_{n,\widehat{h}_{n}}$ are summarized in the following theorem and proposition. To simplify the proof, we consider the non-random hyperparameters $c_{1},...,c_{K}$ such that when $n_{b}$ is given, $h_{1},...,h_{K}$ are also given. The proof is similar but more tedious when considering $\mathcal{P}_{0}$ as the hyperparameter, i.e., $c_{1},...,c_{K}$ are random.
@@ 4.2 Asymptotic Analyses of the Cor-CFD Estimator
$\mathbb{E}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\alpha^{\prime}(\theta_{0})+\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D\Lambda n^{-1/3}+o\left(n^{-1/3}\right),$ (33)
$\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}(q-K)n^{-2/3}+o\left(n^{-2/3}\right),$ (34)
$\mathrm{MSE}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $3\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3}+\left\{\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{n_{b}}{n}}D\Lambda\left(\sqrt{\frac{n_{b}}{n}}D\Lambda+{\rm{sign}}(B)\sigma(\theta_{0})\right)\right.$
$+\left.\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}(q-K)\right\}n^{-2/3}+o\left(n^{-2/3}\right),$ (35)
where ${\bm{c}}=[|c_{1}|,...,|c_{K}|]^{\top}$, ${\bm{c}^{4}}=\left[c_{1}^{4},...,c_{K}^{4}\right]^{\top}$, $\Lambda=\bm{c^{\top}Pc^{4}}$, ${\bm{P}}=\bm{I}-\bm{X}_{e}(\bm{X}_{e}^{\top}\bm{X}_{e})^{-1}\bm{X}_{e}^{\top}$, $\bm{X}_{e}$ is defined in (15), $q=\left\|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right\|_{2}^{2}$, ${\rm{Diag}}(\bm{c^{-1}})={\rm{Diag}}\left(1/|c_{1}|,...,1/|c_{K}|\right)$, ${\rm{sign}}(B)$ is the sign of $B$, and $||{\bm{v}}||_{2}=\left(\sum_{k=1}^{K}v_{k}^{2}\right)^{1/2}$ for any ${\bm{v}}\in\mathbb{R}^{K}$.
The proof of Theorem 4 is provided in Appendix B.2. By comparing (33), (34) and (35) with (9), it becomes attractive that if $B$ and $D\Lambda$ have opposite signs, the squared bias of $\widehat{\Delta}_{n,\widehat{h}_{n}}$ will be smaller than that of the traditional optimal CFD estimator (see (9)). In addition, when $c_{1},...,c_{K}$ are chosen such that $q\leq K$, the variance will be reduced. Proposition 5, a corollary of Theorem 4, further elaborates on these aspects.
@@
      If $-1-{\rm{sign}}(B)\leq\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}D\Lambda\leq 1-{\rm{sign}}(B)$, then
@@
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|\leq\left(1+\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}\left|D\Lambda\right|\right)\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right).$
@@
These results in Proposition 5 can be easily derived by the same technique as in Theorem 4 and the proof is provided in Appendix B.4. Proposition 5 guides the following conclusions:
@@
    The first result indicates that the Cor-CFD estimator regains the traditional optimal CFD estimator when $n_{b}=o(n)$ (refer to (9) and (36)). This aligns with intuition since we allocate a majority of the samples to gradient estimation rather than constant estimation.
@@
    The second result suggests that if $n_{b}=O(n)$, then the absolute value of the bias of the Cor-CFD estimator converges at the same rate as that of the traditional optimal CFD estimator. In particular, 2.i in Proposition 5 demonstrates that the former is smaller than the latter. On the other hand, in Appendix B.3, we show that $\Lambda$ is typically controllable. Therefore, 2.ii in Proposition 5 illustrates that even if the absolute value of the bias of the Cor-CFD estimator is not reduced, its increment is moderate and will not blow up.
@@
    The third result demonstrates that if $n_{b}=O(n)$ and $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$, the Cor-CFD estimator has a reduced variance compared to the traditional optimal CFD estimator. It is worth mentioning that the condition $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ is a highly sufficient condition of $q\leq K$, and the latter yields a reduced variance (see Theorem 4).
@@
    An appealing observation from Proposition 5 is the robustness of the Cor-CFD estimator w.r.t. $r$, in the sense that it exhibits favorable theoretical performance for $0\leq r\leq 1$. In particular, when $r=1$, i.e., the entire simulation budget is used as the pilot samples, the Cor-CFD estimator remains efficient. This theoretical insight serves as a guide, suggesting that when faced with limited simulation resources, allocating the entire budget to generate pilot samples, and the correlation induced samples can lead to an efficient FD estimator.
The variance reduction and possible bias reduction stems from the correlated property of $\Delta_{j}^{c}(h_{k})$ ($j=1,...,n_{b}$) induced by the substitution of $\mathbb{E}[\Delta(h_{k})],k=1,...,K$, with their regression estimators. To illustrate this, let us impractically assume that $\mathbb{E}\Delta(h_{k}),k=1,...,K$ are all known precisely and remain them in (29) but substituting other unknown variables with their estimators in Table 2. This results in a new transformed estimator of (29), denoted by
$\Delta_{j}^{t}(h_{k})=\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}.$
Replacing $\Delta_{j}^{cor}(h_{k})$ in (32) by $\Delta_{j}^{t}(h_{k})$, leads to a new Cor-CFD estimator:
Based on this estimator, $\widehat{\Delta}_{n,\widehat{h}_{n}}$ can be rewritten as
$\widehat{\Delta}_{n,\widehat{h}_{n}}=\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}-\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}-\mathbb{E}\Delta(h_{k})\right).$ (38)
In Appendix B.5, we show that $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$ has the same asymptotic bias and variance as the traditional optimal CFD estimator. Therefore, by comparing results in Theorem 4 and Appendix B.5, the reasons for variance reduction and potential bias reduction become evident: